OpenStudy (anonymous):

solve the equation. check the solution. -2/x+4=4/x+3

OpenStudy (abmon98):

-2/x-4/x=-1, -6/x=-1, -x=-6, x=6

OpenStudy (anonymous):

uhh.. that's not one of them

OpenStudy (anonymous):

x=8 is that an answer choice

OpenStudy (anonymous):

nope.

OpenStudy (anonymous):

they're all fraction but one